OpenStudy (anonymous):

An object is launched at 19.6 m/s from a height of 58.8 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t^2 +19.6t + 58.8. What is the object's maximum height? A. 58.8 B. 194.04 C. 78.4 D. 117.6

OpenStudy (anonymous):

You are trying to find the maximum of the function, so take the first derivative of the function h(t), solve for t where \[ \frac{d}{dt} h(t) = 0 \], then substitute t into the function h(t) to get the answer. Alternatively you can graph the function h(t) out and find the maximum. This is easily done with a graphing calculator.

OpenStudy (anonymous):

I still can't figuer it out, can you show me how please to set up the equation.

OpenStudy (anonymous):

\[ h(t) = -4.9t^2 +19.6t + 58.8\] \[ \frac{d}{dt} h(t) = h'(t) = -9.8t+19.6 \] \[ h'(t) = 0 \rightarrow 9.8t = 19.6 \] \[ t = \frac{19.6}{9.8} \] \[ h_{max} = h(\frac{19.6}{9.8})\]

OpenStudy (anonymous):

Can you tell me if I am correct in this please? hmax = 58.8(2) = 117.6 meters is max height

OpenStudy (anonymous):

answer is 78.4 meters